# Cislunar Constellation Design for Space Situational Awareness with Time-Expanded Facility Location Problem

### v1

@@ 2.2 Libration Point Orbit
Multiple known families of LPOs exist in the vicinity of L1 and L2; among the members of each family, a discrete subset of LPOs that have integer ratio period with respect to the synodic period $T_{\mathrm{syn}}$ is considered. Following convention, we denote an $M$:$N$ LPO as an orbit with period $T=(N/M)T_{\mathrm{syn}}$. This is motivated by previously reported results that placing observer spacecraft on certain $M$:$N$ resonant LPOs with appropriate phasing is particularly beneficial for long-term SSA activity, as favorable illumination conditions become periodic [2, 13]. We note that this choice is reminiscent of opting for designing Earth observation constellations with repeating Low-Earth orbits that exhibit repeating ground tracks [23].
This work considers the synodic-resonant LPOs in the L1/L2 Lyapunov (L1/L2Ly), L2 Halo (L2HS/L2HN), Butterfly (BtfS/BtfN), Distant Prograde Orbit (DPO), and Distant Retrograde Orbit (DRO) families, with synodic resonances 1:1, 3:2, 2:1, 9:4, 5:2, 3:1, 4:1, and 9:2. This results in 40 LPOs, as shown in Figure 1. The parameters of these LPOs are given in Table 9 in Appendix 7.1. The choice of which resonant LPOs to include as candidate orbits matters if the relative positions of member satellites within a constellation must be repeated over a $k$-multiple of the synodic month, where $k$ is the smallest integer such that $kM/N$ is an integer for all considered $M$:$N$ ratios. This translates to whether the sought observer location and pointing schedule are only required to be transient, or if a repeating, steady-state solution is sought. With our choice of resonant LPOs, the constellation returns exactly to its initial configuration every 4 synodic months.
@@
When considering the placement of observers on an LPO, its operational cost should also be taken into account. In this work, we make use of the linear stability index of the LPO to model the station-keeping cost of operating an observer spacecraft on a given LPO. The linear stability index is given by
@@
where $\lambda_{\max}=\max\operatorname{spec}\{\Phi(T,0)\}$.
@@ 3 Demand and Visibility Definitions
In this Section, we introduce the demands considered for the CSSA problem. Specifically, we first distinguish between static and dynamic demands; the former consists of coverage requirement that spans the same volume over time, while the latter consists of time-dependent coverage, either to keep custody of known targets or to monitor transfer corridors that are time-dependent. In both cases, demands are defined by considering a finite set of targets distributed in space; in the static case, the demand for all targets is assumed to be always active, while in the dynamic case, the demand is activated only at time steps when the target exists. Consider $k=1,\ldots,q$ targets and $t=1,\ldots,\ell$ time steps; we define the demand matrix $D\in\mathbb{B}^{\ell\times q}$ as
@@
(a) Static target in lunar sphere of influence
(b) Static target in cone of shame
(c) Dynamic target in low-energy transfer window
Figure 2: Cislunar observation targets for persistent monitoring, shown in Moon-centered Earth-Moon rotating frame centered at the Moon.
3.1 Static Demand
A static demand is defined by a set of targets distributed in space that remain stationary. While in reality, an object cannot remain at a fixed location in space, each “target” in this scenario may be understood as placeholders for defining a volume of interest. Then, the coverage of such a static demand translates to perpetual coverage of the volume that is spanned by the spatial distribution of these static targets.
3.1.1 Persistent Coverage of Lunar Sphere of Influence
@@
3.1.2 Persistent Coverage of Cone of Shame
The so-called cone of shame has been coined by the Air Force Research Laboratory (AFRL) to designate the lunar exclusion zone, within which tracking objects with Earth-based sensors is particularly challenging. In this work, we consider a set of target points distributed in a cone between twice the GEO altitude and Earth-Moon L2, with a cone angle of $30^{\circ}$, as shown in Figure 2(b), consisting of 304 targets.
3.2 Dynamic Demand
3.2.1 Persistent Coverage of Low-Energy Transfer Transit Window
[…]
Figure 3: 200 sample LETs from pre-computed database in the Earth-Moon rotating frame centered at the Earth-Moon system barycenter. Red markers indicate transit window monitoring targets.
Figure 4: Target sparsity pattern of demand matrix $D\in\mathbb{B}^{\ell\times q}$ for persistent coverage of low-energy transfer arrival window over a span of 120 time-steps lasting for four synodic months
@@ 3.3.1 Sun Illumination Model
The Sun illumination is modeled assuming the Earth-Moon orbital plane and Sun-Earth orbital plane to be coplanar. This assumption makes the Sun, Earth, and Moon’s relative geometry periodic, thus allowing for analyses to be generalized for any epoch. Note that the $\sim 5^{\circ}$ difference between the two planes makes a relatively small impact on the illumination conditions, thus justifying this assumption. Let $m_{S}=-26.74$ be the apparent magnitude of the Sun, $d$ be the target diameter, $a_{\mathrm{spec}}$ be its specular reflectance, and $a_{\mathrm{diff}}$ be its diffusion reflectance. Furthermore, let $\boldsymbol{r}_{k}$ denote the position vector of a target, and $\boldsymbol{r}_{j}$ denote the position vector of an observer. Adopting the model from Vendl and Holzinger [2], the apparent magnitude of the target $\bar{m}_{\mathrm{target}}(\boldsymbol{r}_{j},\boldsymbol{r}_{k})$ is given by
$\bar{m}_{\mathrm{target}}(\boldsymbol{r}_{j},\boldsymbol{r}_{k})=m_{\mathrm{S}}-2.5\log_{10}\left(\dfrac{d^{2}}{\|\boldsymbol{r}_{k}-\boldsymbol{r}_{j}\|^{2}}\left[\frac{a_{\mathrm{spec}}}{4}+a_{\mathrm{diff}}p_{\mathrm{diff}}(\phi_{S})\right]\right)$ (6)
where $p_{\mathrm{diff}}(\phi_{S})$ is the diffuse phase angle function given by
$p_{\mathrm{diff}}(\phi_{S})=\frac{2}{3\pi}[\sin(\phi_{S})+(\pi_{S}-\phi_{S})\cos(\phi_{S})]$ (7)
and $\phi_{S}$ is the solar phase angle given by
@@
where $\boldsymbol{l}_{jk}$ is the unit line-of-sight vector from slot $j$ to target $k$, and $\boldsymbol{l}_{Sk}$ is the line-of-sight vector from the Sun to target $k$. Figure 5 shows the contour of $\bar{m}_{\mathrm{target}}$ centered at the observer, with $d=1\,\mathrm{m}$, $a_{\rm spec}=0.0$ and $a_{\rm diff}=0.2$; the contour demonstrates the strong variation of $\bar{m}_{\mathrm{target}}$ with respect to $\phi_{S}$, and the weaker variation with respect to the distance $\|\boldsymbol{r}_{k}-\boldsymbol{r}_{j}\|$ as the latter increases.
@@ 3.3.2 Pointing Direction
Instead of allocating observers to look at a specific target, the TE-$p$-MP will optimize for allocations of observers to a particular pointing direction; whether a target is visible to an observer will then depend on whether the target is within the observer’s field of view. We consider a discrete set of $m$ pointing directions, indexed by $i=1,\ldots,m$, which divides the sphere surrounding an observer into $m$ directions, parameterized in terms of azimuth and elevation angles, $\alpha$ and $\beta$.
The problem of dividing a sphere into $m$ equally distributed points is non-trivial, and suboptimal, approximation approaches such as the Fibonacci spiral may be used, especially for large numbers of $m$. In this work, because the TE-$p$-MP aims to rapidly explore the impact of different parameters of the SSA architecture, we choose to limit $m$ to a relatively small number with easily interpreted directions. Thus, we define a set of 14 pointing directions, with 6 orthogonal directions, and 8 diagonal directions, as shown in Figure 6.
@@ 3.3.3 Visibility boolean
In this work, the emphasis is placed on whether an object is detectable at all by the sensor, rather than the optical observation quality of the target. To this end, using the expression for the apparent magnitude (6), we compute the visibility boolean $\bar{m}^{b}\in\mathrm{B}$, which denotes whether a target $k$ is visible to an observer location $j$ at a given time $t$. Three conditions must be met for an observer to be visible: firstly, the Moon must not be in occultation; secondly, the target must be in the FOV of the observer; lastly, the apparent magnitude $\bar{m}_{\mathrm{target}}(\boldsymbol{r}_{j},\boldsymbol{r}_{k})$ must be below a threshold apparent magnitude $\bar{m}_{\mathrm{crit}}$. Thus, for a given observer position $\boldsymbol{r}_{j}(t)$, target position $\boldsymbol{r}_{k}(t)$, and sensor direction $i(t)$, $\bar{m}^{b}$ is given by
$\bar{m}^{b}(\boldsymbol{r}_{j}(t),\boldsymbol{r}_{k}(t),i(t))$ =$=\begin{cases}0&\cos^{-1}\left(\boldsymbol{l}_{jM}(t)\cdot\boldsymbol{l}_{jk}(t)\right)<\phi_{\mathrm{Moon}}^{\mathrm{crit}}\\ 0&\boldsymbol{r}_{k}(t)\in\operatorname{FOVCone}_{i}\left(\boldsymbol{r}_{j}(t)\right)\wedge\\ 0&\bar{m}_{\mathrm{target}}>\bar{m}_{\mathrm{crit}}\\ 1&\text{otherwise}\end{cases}$ (9)
where $\boldsymbol{l}_{jM}$ is the line-of-sight vector from slot $j$ to the Moon. In equation (9), the second condition ensures the target $\boldsymbol{r}_{k}$ is within the cone formed by the FOV from $\boldsymbol{r}_{j}$ pointed direction $i$. Figure 7 shows the definitions of the quantities defined to compute the visibility boolean.
@@ 4 Time-Expanded $p$-Median Problem
The traditional $p$-Median problem considers the problem of choosing $p$ facilities from $n$ candidate slots to allocate services to $m$ clients. While the mathematical formulation developed in this work shares some resemblance to the terrestrial problem, we note that some of the physical representations of variables are different.
@@
In the remainder of this Section, we begin by introducing the design variables involved in the TE-$p$-MP. Then, we introduce the TE-$p$-MP, along with its associated coefficients, are presented.
@@ 4.1 Design Variables
As in the traditional FLPs, there are two sets of binary variables $X\in\mathbb{B}^{m\times n\times\ell}$ and $Y\in\mathbb{B}^{n}$. The variable $Y$ dictates whether a candidate facility location is to be used, while the variable $X$ dictates the allocation of existing facilities to clients. In the context of the CSSA problem, this allocation can be understood as an assignment of a facility to observe in a particular direction; furthermore, due to the dynamic nature of space, allocation decisions must be taken at each time step. The variables $Y$ and $X$ thus correspond to
$Y_{j}$ =$=\begin{cases}1&\text{observer location $j$ is used}\\ 0&\text{otherwise}\end{cases}$ (10)

### v2

@@ 2.3.2 Libration Point Orbit
Multiple known families of LPOs exist in the vicinity of L1 and L2; among the members of each family, a discrete subset of LPOs have periods approximately equal to an integer ratio of the synodic period $T_{\mathrm{syn}}$, the duration between two consecutive Sun-Earth-Moon alignment. We denote an $M$:$N$ LPO as an orbit with period $T=(N/M)T_{\mathrm{syn}}$. In this work, we focus on a subset of such $M$:$N$ resonant LPOs as part of the candidate observer orbits, following previous works reporting their favorable illumination conditions for CSSA applications [2, 13]. We note that this choice is reminiscent of opting for designing Earth observation constellations with repeating Low-Earth orbits that exhibit repeating ground tracks [57]. An actual spacecraft deployed in a cislunar constellation will be subject to full-ephemeris dynamics subject to various perturbations. Furthermore, due to the Moon’s and Earth’s eccentricities, $T_{\mathrm{syn}}$ is not constant. Thus, in reality, the $M$:$N$ resonance is only an approximation. Nevertheless, we assume the spacecraft will be actively controlled via occasional station-keeping maneuvers to remain in the approximate vicinity of the quasi-periodic LPO that geometrically resembles the CR3BP LPO. In this work, we thus assume the illumination conditions after $M$ revolutions for a given $M$:$N$ resonant LPO are perfectly repeating. The interested reader is pointed to Klonowski et al. [25] for an analysis of performance variation due to this drift.
This work considers the synodic-resonant LPOs in the L1/L2 Lyapunov, Southern and Northern L2 Halo, Southern and Northern Butterfly, Distant Prograde Orbit (DPO), and Distant Retrograde Orbit (DRO) families, with synodic resonances 1:1, 3:2, 2:1, 9:4, 5:2, 3:1, 4:1, and 9:2. This selection results in 40 LPOs, as shown in Figure 1. The parameters of these LPOs are given in Table 7 in Appendix 7.1. Each LPO is constructed via a fixed-time variant of the single-shooting differential correction approach [58], using the LPO entry in the JPL Three-Body Periodic Orbit Catalog [^4] https://ssd.jpl.nasa.gov/tools/periodic_orbits.html with the closest period to the sought resonant period as initial guess. The choice of which resonant LPOs to include as candidate orbits matters if the relative positions of member satellites within a constellation must be repeated over a $N_{\rm syn}$-multiple of the synodic month, where $N_{\rm syn}$ is the smallest integer such that the product $N_{\rm syn}M/N$ is an integer for all considered $M$:$N$ ratios. This translates to whether the sought observer location and pointing schedule are only required to be transient, or if a repeating, steady-state solution is sought. With our choice of resonant LPOs, the constellation returns exactly to its initial configuration every 4 synodic months. We note that the TE-$p$-MP formulation developed in later Sections does not explicitly depend on the $N_{\rm syn}$-month repeating pattern of candidate LPOs, and thus one may instead choose to use an arbitrary set of candidate observer LPOs. However, the number of discretized time steps necessitated to solve for a CSSA constellation and sensor-tasking over an extended duration, for example, lasting multiple years, can be reduced by several folds.
@@
When designing a constellation on LPOs, each member spacecraft’s operational cost should also be considered. The operational cost is a nontrivial quantity that may include various aspects such as station-keeping cost, navigation requirements, and uplink/downlink frequency; the development of a unifying metric is beyond the scope of this work. Instead, assuming the primary driving cost comes from station-keeping activities, the linear stability index $\nu$, defined by
@@
where $\lambda_{\max}$ is the largest eigenvalue of $\Phi(T,0)$, provides insight into the station-keeping maneuver $\Delta V$; Folta et al. [59] demonstrated a monotonic relationship between $\nu$ and $\Delta V$ on a per-LPO-family basis. In the context of this work, the operational cost merely serves the purpose of a secondary metric, used only if two different constellations provide the same coverage capability; thus, using $\nu$ as a proxy to represent operational cost is deemed sufficient.
@@ 3 Demand and Visibility Definitions
We introduce the demands considered for the CSSA problem. Specifically, we first distinguish between static and dynamic demands; the former consists of coverage requirement that spans the same volume over time, while the latter consists of time-dependent coverage, either to keep custody of known targets or to monitor transfer corridors that are time-dependent. In both cases, demands are defined by considering a finite set of targets distributed in space; in the static case, the demand for all targets is assumed to be always active, while in the dynamic case, the demand is activated only at time steps when the target exists. Consider $k=1,\ldots,q$ targets and $t=1,\ldots,\ell$ time steps; we define the demand matrix $D\in\mathbb{B}^{\ell\times q}$ as
@@
(a) Static target in cone of shame
(b) Dynamic target in low-energy transfer window
Figure 2: Cislunar observation targets for persistent monitoring, shown in Earth-Moon rotating frame centered at the Earth-Moon barycenter. Red shaded volume represents the region of interest approximated by the discrete set of targets.
3.1 Static Demand: Cone of Shame
A static demand is defined by a set of targets distributed in space that remain stationary. While in reality, an object cannot remain at a fixed location in space, each “target” in this scenario may be understood as placeholder for defining a volume of interest. Then, the coverage of such a static demand translates to perpetual coverage of the volume that is spanned by the spatial distribution of these static targets.
The static demand considered in this work is the so-called cone of shame, coined by the Air Force Research Laboratory (AFRL) to designate the lunar exclusion zone, within which tracking objects with Earth-based sensors is particularly challenging [60]. In this work, we consider a set of target points distributed in a cone between twice the GEO altitude and Earth-Moon L2, with a cone angle of $30^{\circ}$, as shown in Figure 2(a), consisting of 304 targets.
3.2 Dynamic Demand: Low-Energy Transfer Transit Window
@@
The volume of interest is represented by an orthogonal grid of 675 targets, as shown in Figure 2(b). Since LETs leverage the Sun’s tidal effects, they are inherently tied to a specific time of the month when they cross the neck region and arrive in the lunar vicinity. Thus, while this orthogonal grid remains static in space, persistent coverage of LETs can be achieved by providing coverage at each time step to portions of the grid where at least one LET transits through. With a pre-computed database of LETs developed by the authors [30], we define $D_{tk}$ such that the existence of the target $k$ at a given time $t$ is based on the existence of a LET that passes near target $k$ at this given time. Figure 4 shows the sparsity pattern of $D$ over four synodic periods, which repeats every synodic month, as indicated by the red lines marking the end of each month.
Figure 3: Subset of 1000 sample LETs from pre-computed database in the Earth-Moon rotating frame centered at the Earth-Moon system barycenter. Grey markers indicate transit window monitoring targets.
Figure 4: Target sparsity pattern of demand matrix $D\in\mathbb{B}^{\ell\times q}$ for persistent coverage of low-energy transfer arrival window over a span of 120 time steps lasting for four synodic months
@@ 3.3.1 Sun Illumination Model
Let $m_{S}=-26.74$ denote the apparent magnitude of the Sun, $\boldsymbol{r}_{k}$ denote the position vector of a target, $\boldsymbol{r}_{j}$ denote the position vector of an observer, and let $\phi_{S}$ denote the solar phase angle given by
@@
where $\boldsymbol{l}_{jk}$ is the unit line-of-sight vector from slot $j$ to target $k$, and $\boldsymbol{l}_{Sk}$ is the line-of-sight vector from the Sun to target $k$. The apparent magnitude of a target object is given by [64]
$\bar{m}_{\mathrm{target}}(\boldsymbol{r}_{j},\boldsymbol{r}_{k})=m_{S}-2.5\log_{10}\left[\dfrac{R_{\rm targ}^{2}}{\|\boldsymbol{r}_{k}-\boldsymbol{r}_{j}\|^{2}}\left(C_{\rm diff}p_{\rm diff}(\phi_{S})+\dfrac{1}{4}C_{\rm spec}\right)\right]$ (17)
where $C_{\rm diff}$ is the diffuse reflectance coefficient, $C_{\rm spec}$ is the diffuse reflectance coefficient, $R_{\rm targ}$ is the target’s radius, and $p_{\rm diff}(\phi_{S})$ is the diffuse phase function given by
$p_{\rm diff}(\phi_{S})=\dfrac{2}{3\pi}[\sin(\phi_{S})+(\pi-\phi_{S})\cos(\phi_{S})]$ (18)
In this work, we consider targets with $R_{\rm targ}=2\,\mathrm{m}$, $C_{\rm diff}=0.2$, and $C_{\rm spec}=0$. Figure 5 shows the contour of $\bar{m}_{\mathrm{target}}$ centered at the observer with $\phi_{S}=200^{\circ}$; the contour demonstrates the strong variation of $\bar{m}_{\mathrm{target}}$ with respect to $\phi_{S}$, and the weaker variation with respect to the distance $\|\boldsymbol{r}_{k}-\boldsymbol{r}_{j}\|$ as the latter increases.
@@
In this work, the line of sight to the Sun $\boldsymbol{l}_{Sk}$ is assumed to lie on the Earth-Moon orbital plane, rotating clockwise at a constant angular velocity $\omega_{S}$ about the $z$-axis of the Earth-Moon rotating frame. This assumption is based on the small, $\sim 5^{\circ}$ difference between the Sun-Earth and Earth-Moon orbital plane, making a relatively small impact on the illumination conditions. Adopting this assumption, the illumination condition is exactly one synodic month-periodic, thereby making a given CSSA architecture and the corresponding sensor-tasking schedule an approximate solution for any epoch. Note, however, that this assumption is not a necessary ingredient to the subsequent development of the TE-$p$-MP; if one opts to incorporate $\boldsymbol{l}_{Sk}$ with $z$ components, the TE-$p$-MP solution becomes epoch-specific.
@@ 3.3.2 Pointing Direction
Instead of allocating observers to look at a specific target, the TE-$p$-MP will optimize for allocations of observers to a particular pointing direction; whether a target is visible to an observer will then depend on whether the target is within the observer’s field of view. We consider a discrete set of $m$ pointing directions, indexed by $i=1,\ldots,m$, which divides the sphere surrounding an observer into $m$ directions, parameterized in terms of azimuth and elevation angles.
The problem of dividing a sphere into $m$ equally distributed points is non-trivial, and suboptimal, approximation approaches such as the Fibonacci spiral may be used, especially for large numbers of $m$. In this work, because the TE-$p$-MP aims to rapidly explore the impact of different parameters of the SSA architecture, we choose to limit $m$ to a relatively small number with easily interpreted directions. Thus, we define a set of 14 pointing directions, with 6 orthogonal directions, and 8 diagonal directions, as shown in Figure 6. We assume an azimuth and elevation angles of of $0^{\circ}$ points the sensor along the $x$-axis in the Earth-Moon rotating frame.
@@ 3.3.3 Visibility boolean
We seek to define a binary parameter that dictates whether an observation of a target is possible and is of sufficient quality. We make use of the target’s apparent magnitude (17) to define the visibility boolean, denoted by $\bar{m}^{b}\in\mathrm{B}$; the use of apparent magnitude as a metric for assessing whether a target is observable can be found in previous literature, for example in [2, 13]. Alternatively, one may also incorporate more advanced observation metrics such as the signal-to-noise ratio (SNR) [15, 1, 9, 25] to define the visibility boolean; since the focus of this work is to propose the algorithmic framework for formulating and solving the CSSA constellation design problem, the incorporation of such metrics are deemed beyond the scope of this work. Four conditions must be met for an observer to be visible: the first and second conditions ensure the Moon and the Earth are not ahead or behind the target; the third condition ensures the target to lie in the FOV of the observer; the last condition ensures the target’s apparent magnitude $\bar{m}_{\mathrm{target}}(\boldsymbol{r}_{j},\boldsymbol{r}_{k})$ to be below a threshold visible apparent magnitude $\bar{m}_{\mathrm{crit}}$. Taking these conditions into account, for a given observer position $\boldsymbol{r}_{j}(t)$, target position $\boldsymbol{r}_{k}(t)$, and sensor direction $i(t)$, $\bar{m}^{b}$ is given by
$\bar{m}^{b}(\boldsymbol{r}_{j}(t),\boldsymbol{r}_{k}(t),i(t))$ =$=\begin{cases}0&\cos^{-1}\left(\boldsymbol{l}_{jM}(t)\cdot\boldsymbol{l}_{jk}(t)\right)<\phi_{\mathrm{Moon}}^{\mathrm{crit}}(\boldsymbol{r}_{j})\\ 0&\cos^{-1}\left(\boldsymbol{l}_{jE}(t)\cdot\boldsymbol{l}_{jk}(t)\right)<\phi_{\mathrm{Earth}}^{\mathrm{crit}}(\boldsymbol{r}_{j})\\ 0&\boldsymbol{r}_{k}(t)\notin\operatorname{FOVCone}_{i}\left(\boldsymbol{r}_{j}(t)\right)\\ 0&\bar{m}_{\mathrm{target}}>\bar{m}_{\mathrm{crit}}\\ 1&\text{otherwise}\end{cases}$ (19)
where $\boldsymbol{l}_{jk}$ is the line of sight vector from slot $j$ to target $k$, $\boldsymbol{l}_{jM}$ is the line-of-sight vector from slot $j$ to the Moon, and $\boldsymbol{l}_{jE}$ is the line-of-sight vector from slot $j$ to the Earth. In equation (19), the first condition ensures the angular separation between $\boldsymbol{l}_{jk}$ and $\boldsymbol{l}_{jM}$ is separated by the lunar apparent angular radius $\phi_{\mathrm{Moon}}^{\mathrm{crit}}(\boldsymbol{r}_{j})$; similarly, the second condition ensures the angular separation between $\boldsymbol{l}_{jk}$ and $\boldsymbol{l}_{jE}$ is separated by the Earth’s apparent angular radius $\phi_{\mathrm{Earth}}^{\mathrm{crit}}(\boldsymbol{r}_{j})$; the third condition ensures the target $\boldsymbol{r}_{k}$ is within the cone formed by the FOV from $\boldsymbol{r}_{j}$ pointed direction $i$; finally, the fourth condition ensures the target appears sufficiently bright to the observer. Figure 7 shows the definitions of the quantities defined to compute the visibility boolean.
@@ 4 Time-Expanded $p$-Median Problem
The traditional $p$-Median problem considers the problem of choosing $p$ facilities from $n$ candidate slots to allocate services to $m$ clients. While the mathematical formulation developed in this work shares some resemblance to the problem in terrestrial applications, we note that some of the physical representations of variables are different.
@@ 4.1 Design Variables
The TE-$p$-MP has two sets of binary variables, $X\in\mathbb{B}^{m\times n\times\ell}$ and $Y\in\mathbb{B}^{n}$. In traditional FLPs, the variable $Y$ dictates whether a candidate facility location is used in the solution, while the variable $X$ dictates the allocation of existing facilities to clients. In the context of the CSSA problem, this allocation can be understood as an assignment of an observer to monitor in a particular direction; furthermore, due to the dynamic nature of space, allocation decisions must be taken at each time step. The variables $Y$ takes the same definition as (1), while variables $X$ correspond to